Use the Cauchy-Schwarz inequality to solve x³-25x²- 4x+ 100= 0 if we know that all its roots are rational.
1
Expert's answer
2020-02-28T10:12:01-0500

"x^{3}-25x^{2}-4x+100=0" then "x^{2}(x-25)-4(x-25)=0" then "(x^{2}-4)(x-25)=0" hence "x_1=2,x_2=-2,x_3=25"